3 wholes and 2 4th x1 wholes 3 twelveths
Answer by richwmiller(17219)

(3x+2/4)(1x+3/12)
3 2/4*1 1/4=
(14/4)*(5/4)=7/2*5/4=35/8=4 3/8
or is tweths really tenths?
3.5*1.25=4.375= 4 and 3/8